### v1

@@ 4 Methodology
For the first point, our method defines the significant bi-cliques to be those corresponding to concepts in the core iceberg concept lattice. These significant bi-cliques can be extracted without exhaustive enumeration and are expected to be indeed significant for link prediction. For the second point, we have developed a modified Transformer encoder network called unordered Transformer encoder to effectively capture the information and make link predictions. We will first give a brief introduction to these techniques we develop before giving the detailed working flow of our method.
@@ 4.1 Core iceberg concept lattice
In our method, we define a concept $(A,B)$ to be significant if $l_{1}\leq\|A\|\leq u_{1}$ and $l_{2}\leq\|B\|\leq u_{2}$. That is, a concept is significant if both its intents’ and extents’ lengths are neither too long nor too short. According to the definition of an iceberg concept lattice, we can find that:
  • All concepts $(A,B)$ such that $|A|\geq l_{1}$ and $|B|\leq u_{2}$ form into an “upward” iceberg concept lattice.
  • All concepts $(A,B)$ such that $|A|\leq u_{1}$ and $|B|\geq l_{2}$ form into a “downward” iceberg concept lattice.
@@
Since the concepts in the iceberg concept lattice can be extracted without exhaustive enumeration by algorithms like LCM [30], by applying a simple modification to the LCM algorithm, we will be able to extract the significant concepts in the core iceberg concept lattice without exhaustive enumeration. Details will be introduced in the workflow of our method.
Certainly, if the significant concepts are only easy to extract, but are not really significant in terms of the information they can provide for link prediction, extracting them would be meaningless. However, we consider the concepts in the core iceberg concept lattice to be indeed significant for link prediction because they encompass objects and attributes that frequently appear in concepts [18].
@@ 4.2 Unordered Transformer Encoder
Previous research has utilized many different methods for capturing the information from formal concepts, including pure rule-based method [12], matrix-factorization-based methods [14], and deep-learning-based methods [27, 16, 17]. While in this research, we consider the Transformer encoder most suitable for the following reason. Formal concepts can be regarded as unordered sequences. The most valuable information a formal concept can provide is the closed relations between objects and attributes within its extent and intent [20]. The Transformer network is designed for extracting dependencies between elements in a sequence, so we believe it can capture the information provided by a formal concept well.
As mentioned in the previous section, we consider an encoder-only network more suitable for this task, because the decoder works recursively and thus it strongly relies on the order of the elements in the sequence, while formal concepts are completely unordered sequences. Furthermore, the original version of the Transformer encoder still has a mechanism called positional embedding, which memorizes the order of elements in the input sequence. Since the extents and intents are completely unordered sequences, in our method, we have removed this positional encoding mechanism. We name our modified Transformer architecture an unordered Transformer encoder and believe it is well suitable for our task.
@@ 4.3 The workflow of our method
Significant Concept Extraction: In this step, we convert the bipartite network into a formal context. Then, we set the upper and lower bounds on the lengths of extents and intents, i.e., $u_{1},u_{2},l_{1}$, and $l_{2}$, and only extract concepts in which the lengths of extents and intents fall in this range. That is, we are to search all significant concepts in the core iceberg concept lattice. Such a searching process is achieved by an algorithm derived from the LCM algorithm [30]. Please refer to Appendix A for a detailed algorithm.
The upper and lower bounds for the lengths of extents and intents should be determined based on the size and density of formal contexts, as well as the available computational resources. If the bounds are too restrictive, there may be no concepts in the core iceberg concept lattice. On the other hand, too loose bounds may result in an excessive number of extracted concepts, which can be computationally expensive. In practice, the choice of bounds involves a trade-off between prediction performance and computational resources.
@@
After generating the four intermediate sets, we can generate the concept samples. The positive concept samples are generated from extents $E_{\mathrm{p}}$ and intents $I_{\mathrm{p}}$, denoted as $C_{\mathrm{p}}$. The negative samples are generated from non-extent object sets $E_{\mathrm{n}}$ and non-intent attribute sets $I_{\mathrm{n}}$, denoted as $C_{\mathrm{n}}$. The details and the corresponding pseudocode are presented in B.2.
@@
After obtaining all training samples, we need to pad them to the same length with special tokens “[PAD]” so that they can be fed into the network. We find the length of the longest sequence and pad all sequences to this length. The details and the corresponding pseudocode are presented in B.4. After this step, we will obtain the the final training set $T_{\mathrm{F}}$, with its label (positive or negative) set.
Training: We employ an unordered Transformer encoder and a Multi-Layer Perceptron (MLP) to train a model for link prediction. The model takes sequences concatenated by object sets and attribute sets as input. A special token [CLS] is added at the first position of an input sequence. After the training step, the final embedding of this [CLS] token is expected to capture and represent the information of the whole sequence. Another special token [SEP] is added between object sets and attribute sets to help the model differentiate two different sets. The detailed structure of the Transformer encoder model is illustrated in Fig 7.
@@
In our experiments, the model contains $N=9$ identical encoder layers, each of which contains a multi-head attention layer with $H=12$ heads and a feedforward neural network. The input size is $(n_{\mathrm{obj}}+n_{\mathrm{attr}})\times(l_{\mathrm{ext}}+l_{\mathrm{int}})$, where $n_{\mathrm{obj}}$ is the number of objects, $n_{\mathrm{attr}}$ is the number of attributes, $l_{\mathrm{ext}}$ is the length of the longest extents, and $l_{\mathrm{int}}$ is the length of the longest intent.
@@
With the network for training setup, we are to feed the training samples into the network to train the model. For each sample $(X,Y)$ from the final training set $T_{\mathrm{F}}$, “$[CLS]+X+[SPE]+Y$ ” is fed into the model. The training process may be conducted multiple times (epochs) until the loss has converged.
4.4 Analysis of our method
@@
First, in contrast to previous FCA-based methods, our method is applicable to large datasets with the help of the core iceberg concept lattice. While previous FCA-based methods require the full concept lattice for link prediction, our method selects significant concepts to avoid the requirement for an exhaustive enumeration of all concepts. Furthermore, even if the significant concepts we select happen to be meaningless for link prediction, our method can still make link predictions from the raw information in the formal context, which further enhances its flexibility and efficiency.
Second, our method offers flexibility in adjusting computational resources according to specific needs. As mentioned above, there is a trade-off between prediction performance and computational resources. Extracting a greater number of concepts typically enhances prediction performance but also increases the time required for concept extraction and model training. Additionally, longer extents and intents will lead to a larger model size, which will require more computational resources during training. One can decide the number and the length of extracted concepts by defining the range of lengths for extents and intents based on the available computational resources. This adaptability ensures efficient resource utilization while maintaining competitive prediction performance.
Additionally, in our method, we propose a Transformer network which is specifically for bipartite link prediction. We remove the “positional embedding” and add special tokens “[CLS]” and “[SEP]” to make the Transformer architecture suitable for tasks in bipartite networks so that it can directly and effectively learn the pairwise relationships between two nodes from different sets in the bipartite networks. With the help of the powerful Transformer architecture, The negative influence on prediction performance caused by the information not contained in the extracted formal concepts may be counterbalanced. Moreover, the proposed network can directly process bipartite networks to make link predictions with high performance. We will demonstrate the performance of it in the ablation experiments.
@@ 5.1 Datasets
HetRec: The HetRec dataset [31] represents a user-movie rating dataset collected from different websites. Originally, the dataset consisted of users rating movies on a scale of 1 to 5. To transform it into a binary dataset, we consider the scores equal to or greater than 3 as indications of user preferences(favorites), while the scores below 3 represent dislikes. This conversion transforms the original dataset into a binary user-movie dataset. In this bipartite network, users represent one set of nodes, movies represent the other set, and an edge denotes a user liking a particular movie. Our objective is to simulate a scenario where we predict movies that users may like in the future, akin to a movie recommendation system. We follow the same methodology used for CDT and BMS-POS datasets, utilizing the original network as the target for prediction and generating the input network by randomly removing $10\%$ of user-movie edges.
MovieLens: This dataset, like HetRec, is also a user-movie rating dataset generated from the MovieLens 25M dataset [32]. We consider the scores equal to 5 as indications of user preferences, while the scores below 5 represent dislikes to transform the original dataset into a binary relational dataset. Then we remove the users who rated fewer than 300 movies and the movies which are rated by fewer than 120 users from the datasets. The objective and the methodology to generate the input and target networks are the same as those used in the HetRec dataset.
@@ 5.2 Experiments
The baseline methods we compare our method with are as follows: several heuristic rule-based link prediction methods like Common Neighbors (CN) [7], Adamic-Adar Coefficient (AA) [13], Jaccard Coefficient (JC) [33], Resource Allocation(RA) [34], Rooted PageRank(RPR) and SimRank(SR) [35]; a classic bipartite link prediction method, Matrix factorization with singular value decomposition(MF-SVD); an embedding-based method, random walk, which trains node embedding using a random walk process and then trains a linear regression classifier on samples of present links with the features of nodes set to their embeddings to make predictions; a novel GNN-based link prediction method,SBGNN, which is specifically designed for bipartite network [36]. The results are reported in Table 2.
@@
From the results, we can observe that BicliqueEncoder has overall better performances than other baseline methods across all five datasets. For MF-SVD, while it exhibits higher AUC scores than BicliqueEncoder across all datasets, its overall performance is still considered lower than BicliqueEncoder because of its low AUPR scores. According to previous research, the AUPR score gives more weight to the positive samples. A low AUPR score indicates that the model’s high prediction scores do not correlate well with being in the positive class, which suggests that the model has difficulty achieving high precision [37]. Since in link prediction we focus on predicting the generation of new links, which is predicting the positive samples but not the negative samples [38, 14], so if a model exhibits a low AUPR, it is considered to have a low performance in link prediction.
@@ 5.3 Ablation Experiment
The results indicate that learning information from formal concepts improves the prediction performance across these datasets, although the extent of improvement varies. Particularly, the improvement on BMS-POS is minimal compared to the other datasets. This may be due to the relatively small number of the extracted concepts used to train the model in comparison to the other datasets. This demonstrates that extracting more concepts and utilizing them to train the model can enhance the prediction performance. Additionally, the varying importance of the information contained in the extracted concepts for link prediction may contribute to the differences in improvement across datasets. In some datasets, the extracted concepts might be more representative of the entire concept lattice and thus more useful for improving link predictions compared to other datasets. Another possible reason is that the maximal length of concepts varies across datasets, resulting in different input lengths for the model. Longer input sequences may make it harder to train the model, which counterbalances the improvement brought by adding concepts as training samples to some extent.
@@ 5.4 Supplementary Experiment
To evaluate our method on small datasets and compare it with the previous FCA-based methods, we apply it to the datasets used in the publication where BERT4FCA was proposed [17]. We apply BicliqueEncoder on these datasets and compare the prediction performance and execution time with BERT4FCA. We choose not to further compare our method with other FCA-based methods on the above five datasets because first, BERT4FCA was proven the best FCA-based link prediction method [17]; second, previous methods are extremely time-consuming and require significant computational resources that is unavailable to us. Since the sizes of these small datasets are manageable, we extract all the concepts and utilize them to make link predictions in BicliqueEncoder. We depict the features of these datasets in Table 4. Please refer to [17] for a detailed description of each dataset. The prediction results are reported in Table 5 and the time consumption is reported in Table 6.
@@ 6 Conclusion and Future Works
In future research, we plan to explore alternative approaches for extracting significant formal concepts from concept lattices, beyond the use of iceberg concept lattices, to enhance the efficiency of FCA-based methods for bipartite link prediction. Additionally, we aim to develop a method that enables the model to learn the order relations of concepts within concept lattices. These order relations, as demonstrated in the BERT4FCA research, are useful for bipartite link prediction. We expect the prediction performance can be further improved by learning this information.
@@ Appendix A The Pseudocode for Significant Concept Extraction
In the “Significant Concept Extraction” step, we set the upper and lower bounds on the lengths of extents and intents and only extract concepts in which the lengths of extents and intents fall in this range. Such a searching process is achieved by an algorithm derived from the LCM algorithm [30]. The pseudocode of the searching algorithm is presented in 1.
@@
Input A formal context $\mathbb{K}=(G,M,I)$; the upper bound for the lengths of extents $u_{1}$; the lower bound for the lengths of the extents $l_{1}$; the upper bound for the lengths of intents $u_{2}$; the lower bound for the lengths of intents $l_{2}$.

### v2

@@ 4 Methodology
For the first point, our method defines the significant bi-cliques to be those corresponding to concepts in the core iceberg concept lattice. These significant bi-cliques can be extracted without exhaustive complete enumeration of all formal concepts and are expected to be indeed significant for link prediction. For the second point, we have developed a modified Transformer encoder network called unordered Transformer encoder to effectively capture the information and make link predictions. We will first give a brief introduction to these techniques we develop before giving the detailed working flow of our method.
@@ 4.1 Core iceberg concept lattice
In our method, we define a concept $(A,B)$ to be significant if $l_{1}\leq\|A\|\leq u_{1}$ and $l_{2}\leq\|B\|\leq u_{2}$. That is, a concept is significant if both its intents’ and extents’ sizes are neither too long nor too short. According to the definition of an iceberg concept lattice, we can find that:
  • All concepts $(A,B)$ such that $|A|\geq l_{1}$ and $|B|\leq u_{2}$ form an “upward” iceberg concept lattice.
  • All concepts $(A,B)$ such that $|A|\leq u_{1}$ and $|B|\geq l_{2}$ form a “downward” iceberg concept lattice.
@@
Since the concepts in the iceberg concept lattice can be extracted without exhaustive enumeration by algorithms like LCM ([32]), which is a frequent closed sets mining algorithm. By applying a simple modification to the LCM algorithm, we will be able to extract the significant concepts in the core iceberg concept lattice without exhaustive enumeration. Details will be introduced in the workflow of our method.
Certainly, if the significant concepts are only easy to extract, but are not really significant in terms of the information they can provide for link prediction, extracting them would be meaningless. However, we consider the concepts in the core iceberg concept lattice to be indeed significant for link prediction because they encompass objects and attributes that frequently appear in concepts, as well as their binary relations ([31]).
@@ 4.2 Unordered Transformer Encoder
Previous research has utilized many different methods for capturing the information from formal concepts, including pure rule-based method ([40]), matrix-factorization-based methods ([25]), and deep-learning-based methods ([7, 22, 26]). While in this research, we consider the Transformer encoder most suitable for the following reason. The most valuable information a formal concept can provide is the closed relations between objects and attributes within its extent and intent ([2]). The Transformer network is designed for extracting dependencies between elements in a sequence, so we believe it can capture the information provided by a formal concept well.
As mentioned in the previous section, we consider an encoder-only network more suitable for this task, because the decoder works recursively and thus it strongly relies on the order of the elements in the sequence, while formal concepts are completely unordered. Furthermore, the original version of the Transformer encoder still has a mechanism called positional embedding, which memorizes the order of elements in the input sequence. Since the extents and intents are completely unordered, in our method, we have removed this positional encoding mechanism. We name our modified Transformer architecture an unordered Transformer encoder and believe it is well suitable for our task.
@@ 4.3 The workflow of our method
Significant Concept Extraction: In this step, we convert the bipartite network into a formal context. Then, we set the upper and lower bounds on the sizes of extents and intents, i.e., $u_{1},u_{2},l_{1}$, and $l_{2}$, and only extract concepts in which the sizes of extents and intents fall in this range. That is, we are to search all significant concepts in the core iceberg concept lattice. Such a searching process is achieved by an algorithm derived from the LCM algorithm ([32]). Please refer to Appendix A for a detailed algorithm.
The upper and lower bounds for the sizes of extents and intents should be determined based on the size and density of formal contexts, as well as the available computational resources. If the bounds are too restrictive, there may be no concepts in the core iceberg concept lattice. On the other hand, too loose bounds may result in an excessive number of extracted concepts, which can be computationally expensive. In practice, the choice of bounds involves a trade-off between prediction performance and computational resources.
@@
After generating the four intermediate sets, we can generate the concept samples. The positive samples consist of all formal concepts, which are pairs of an extent and an intent, denoted as $C_{\mathrm{p}}$. The negative samples are generated from non-extent object sets $E_{\mathrm{n}}$ and non-intent attribute sets $I_{\mathrm{n}}$, denoted as $C_{\mathrm{n}}$. The details and the corresponding pseudocode are presented in B.2.
@@
After obtaining all training samples, we need to pad them to the same length with special tokens “[PAD]” so that they can be fed into the network. We find the length of the longest sequence and pad all sequences to this length. The details and the corresponding pseudocode are presented in B.4. After this step, we will obtain the final training set $T_{\mathrm{F}}$, with its label (positive or negative) set.
Training: We employ an unordered Transformer encoder and a Multi-Layer Perceptron (MLP) to train a model for link prediction. The model takes concatenations of object sets and attribute sets as input. A special token [CLS] is added at the first position of an input sequence. After the training step, the final embedding of this [CLS] token is expected to capture and represent the information of the whole sequence. Another special token [SEP] is added between object sets and attribute sets to help the model differentiate two different sets. The detailed structure of the Transformer encoder model is illustrated in Fig 7.
@@
In our experiments, the model contains $N=9$ identical encoder layers, each of which contains a multi-head attention layer with $H=12$ heads and a feedforward neural network. The input size is $(n_{\mathrm{obj}}+n_{\mathrm{attr}})\times(l_{\mathrm{ext}}+l_{\mathrm{int}})$, where $n_{\mathrm{obj}}$ is the number of objects, $n_{\mathrm{attr}}$ is the number of attributes, $l_{\mathrm{ext}}$ is the size of the longest extents, and $l_{\mathrm{int}}$ is the size of the longest intent.
@@
With the network for training setup, we are to feed the training samples into the network to train the model. For each sample $(X,Y)$ from the final training set $T_{\mathrm{F}}$, “$[CLS]+X+[SEP]+Y$ ” is fed into the model. The training process may be conducted multiple times (epochs) until the loss has converged.
4.4 Discussion of our method
@@
First, in contrast to previous FCA-based methods, our method is applicable to large datasets with the help of the core iceberg concept lattice. While previous FCA-based methods require the full concept lattice for link prediction, our method selects significant concepts to avoid the requirement for an exhaustive enumeration of all concepts. For example, if the number of formal concepts in a formal context is $N$, the time complexity of the FCA concept extraction process in other FCA-based methods is $O(N)$, whereas in our method, it is $O(n)$, where $n$ is significantly smaller than $N$. Furthermore, in extreme cases where the extracted significant concepts provide limited useful information for link prediction, our method can still learn directly from the raw formal context and achieve strong performance, enhancing its flexibility and robustness.
Second, our method offers flexibility in adjusting computational resources according to specific needs. As mentioned above, there is a trade-off between prediction performance and computational resources. Extracting a greater number of concepts typically enhances prediction performance but also increases the time required for concept extraction and model training. Additionally, longer extents and intents will lead to a larger model size, which will require more computational resources during training. One can adjust the number and the length of extracted concepts by defining the range of sizes for extents and intents based on the available computational resources. This adaptability ensures efficient resource utilization while maintaining competitive prediction performance.
Additionally, in our method, we propose a Transformer network which is specifically for bipartite link prediction. We remove the “positional embedding” and add special tokens “[CLS]” and “[SEP]” to make the Transformer architecture suitable for tasks in bipartite networks so that it can directly and effectively learn the pairwise relationships between two nodes from different sets in the bipartite networks. With the help of the powerful attention mechanism in the Transformer encoder networks, the negative influence on prediction performance caused by the information not contained in the extracted formal concepts can be counterbalanced. Moreover, the proposed network can directly process bipartite networks to make link predictions with high performance. We will demonstrate the performance of it in the ablation experiments.
However, our method also has some limitations. Focusing only on significant bi-cliques may result in the loss of minor patterns that are not contained in the extracted bi-cliques. Bi-cliques represent clusters of nodes that are connected to each other. They are important for the model to uncover the patterns of link formation underlying bipartite networks, as they help the model learn to identify pairs of nodes that connect (positive samples). However, nodes that do not connect are also important for link prediction, as they help the model learn to identify pairs of nodes that do not connect (negative samples). Only learning from bi-cliques may not effectively help the model identify negative samples.
The importance of bi-cliques for link prediction also varies across different datasets, leading to varying improvements in prediction accuracy due to the use of bi-cliques. In some datasets, bi-cliques may perfectly reflect the pattern of link formation, resulting in high prediction accuracy, while in others, they may not capture the underlying structure as effectively.
@@ 5.1 Datasets
HetRec: The HetRec dataset ([3]) represents a user-movie rating dataset collected from different websites. Originally, the dataset consisted of users rating movies on a scale of 1 to 5. To transform it into a binary dataset, we consider the scores equal to or greater than 3 as indications of user preferences(favorites), while the scores below 3 represent dislikes. This conversion transforms the original dataset into a binary user-movie dataset. In this bipartite network, users represent one set of nodes, movies represent the other set, and an edge denotes a user liking a particular movie. Our objective is to simulate a scenario where we predict movies that users may like in the future, akin to a movie recommendation system. We follow the same methodology used for CDT and BMS-POS datasets, utilizing the original network as the target for prediction and generating the input network by randomly removing $10\%$ of user-movie edges.
MovieLens: This dataset, like HetRec, is also a user-movie rating dataset generated from the MovieLens 25M dataset ([11]). We consider the scores equal to 5 as indications of user preferences, while the scores below 5 represent dislikes to transform the original dataset into a binary relational dataset. Then we remove the users who rated fewer than 300 movies and the movies which are rated by fewer than 120 users from the datasets. The objective and the methodology to generate the input and target networks are the same as those used in the HetRec dataset.
@@ 5.2 Experiments
The baseline methods we compare our method with are as follows: several heuristic rule-based link prediction methods like Common Neighbors (CN) ([34]), Adamic-Adar Coefficient (AA) ([38]), Jaccard Coefficient (JC) ([17]), Resource Allocation(RA) ([36]), Rooted PageRank(RPR) and SimRank(SR) ([14]); a classic bipartite link prediction method, Matrix factorization with singular value decomposition(MF-SVD); an embedding-based method, random walk, which trains node embedding using a random walk process and then trains a linear regression classifier on samples of present links with the features of nodes set to their embeddings to make predictions; a novel GNN-based link prediction method, SBGNN, which is specifically designed for bipartite network ([12]). The results are reported in Table 2.
@@
From the results, we can observe that BicliqueEncoder has overall better performances than other baseline methods across all five datasets. For MF-SVD, while it exhibits higher AUC scores than BicliqueEncoder across all datasets, its overall performance is still considered lower than BicliqueEncoder because of its low AUPR scores. According to previous research, the AUPR score gives more weight to the positive samples. A low AUPR score indicates that the model’s high prediction scores do not correlate well with being in the positive class, which suggests that the model has difficulty achieving high precision ([6]). Since in link prediction we focus on predicting the generation of new links, which is predicting the positive samples but not the negative samples ([4, 25]), so if a model exhibits a low AUPR, it is considered to have a low performance in link prediction.
@@ 5.3 Ablation Experiment
The results indicate that learning information from formal concepts improves the prediction performance across these datasets, although the extent of improvement varies. The varying importance of the information contained in the extracted concepts for link prediction may contribute to the differences in improvement across datasets. In some datasets, the extracted concepts might be more representative of the entire concept lattice and thus more useful for improving link predictions compared to other datasets. Another possible reason is that the maximal length of concepts varies across datasets, resulting in different input lengths for the model. Longer input sequences may make it harder to train the model, which counterbalances the improvement brought by adding concepts as training samples to some extent.
@@ 5.4 Supplementary Experiment
To evaluate our method on small datasets and compare it with the previous FCA-based methods, we apply it to the datasets used in the publication where BERT4FCA was proposed ([26]). We apply BicliqueEncoder on these datasets and compare the prediction performance and execution time with BERT4FCA. We choose not to further compare our method with other FCA-based methods on the above five datasets because first, BERT4FCA was proven the best FCA-based link prediction method ([26]); second, previous methods are extremely time-consuming and require significant computational resources that is unavailable to us. Since the sizes of these small datasets are manageable, we extract all the concepts and utilize them to make link predictions in BicliqueEncoder. We depict the features of these datasets in Table 4. Please refer to ([26]) for a detailed description of each dataset. The prediction results are reported in Table 5 and the time consumption is reported in Table 6.
@@ 6 Conclusion and Future Works
In future work, we plan to propose a method to determine the optimal number and length of extracted concepts without relying on prior experience or costly trials, thereby improving the practicality of our method. We also intend to explore alternative approaches for extracting significant formal concepts from concept lattices, beyond the use of iceberg concept lattices, to enhance the efficiency of FCA-based methods for bipartite link prediction. Additionally, we aim to extend our method to non-bipartite and heterogeneous networks to broaden its applicability. We also plan to develop a method that enables the model to learn the order relations of concepts within concept lattices. These order relations, as demonstrated in the BERT4FCA research, are useful for bipartite link prediction. We expect the prediction performance can be further improved by learning this information.
@@ Appendix A The Pseudocode for Significant Concept Extraction
In the “Significant Concept Extraction” step, we set the upper and lower bounds on the sizes of extents and intents and only extract concepts in which the sizes of extents and intents fall in this range. Such a searching process is achieved by an algorithm derived from the LCM algorithm ([32]). The pseudocode of the searching algorithm is presented in 1.
@@
Input A formal context $\mathbb{K}=(G,M,I)$; the upper bound for the sizes of extents $u_{1}$; the lower bound for the sizes of the extents $l_{1}$; the upper bound for the sizes of intents $u_{2}$; the lower bound for the sizes of intents $l_{2}$.